Chris Martin worked for myself and Professor James Allwright in a Research Experience for Undergraduates program in computational physics at Syracuse University in the summer of 1993. The project he worked on was an effort to parallelize a program which does Monte Carlo simulation of dynamically triangulated random surfaces, which are discretized models of quantum gravity. The parallel algorithm requires a graph coloring of a triangulated mesh. Chris’s project was to take a number of different sequential graph coloring algorithms and programs and to parallelize them, with an existing parallel graph coloring algorithm as a guide. He was asked to write the programs in two different parallel languages – Express Fortran (a MIMD message-passing language) and CM Fortran (a SIMD data-parallel language) – and to compare their performance on two different parallel computers (an Intel iPSC/860 and a Thinking Machines CM-5) for some sample graphs.

This required an understanding of the graph coloring algorithms, the exisiting code, parallel computing concepts, the two specific parallel languages and computers, and the concepts underlying the physics algorithms and program. In a very short period of time, Chris was able to familiarize himself with all these aspects of the work, and to produce working parallel programs and results with very little supervision. He produced an impressive paper and poster on his work, and he will be co-author on a paper on this work to be submitted to an academic journal.

Chris is very quick to pick up new concepts, is a responsible and enthusiastic worker, and a good programmer. He is very intelligent, amiable, polite, wellspoken, and a pleasure to work with. I believe he would do extremely well in graduate school and recommend him highly.